# Topological BF Theory construction of twisted dihedral quantum double phases from spontaneous symmetry breaking

### v1

@@ Abstract
SI Data of the dihedral quantum double
@@
where the function $\mathrm{Cell}_{n}[x]$ subtracts (or adds) integer multiples of $n$ to take $x$ back to the interval ((n1)/2,(n1)/2](-(n-1)/2,(n-1)/2], and we denoted a generic element of $D_{n}$ by $g=C_{n}^{r}M^{s}$. The $R$-symbol is
@@
SII Equivalence between $S[O(2)\times O(2)]$ BF theory and $D_{n}$ quantum double
SII.1 Ground State Degeneracy and Quantum Dimension
@@
SII.2 Topological Spins and Fusion Rules
@@
Their fusion rule, consistent with those of $\psi$ anyons in $D^{\omega}(D_{n})$, is the same as Eq. (8). To be concrete, here $W_{p,p}$ and $W_{p,\frac{n}{2}-p}$ are identified as $\psi_{l}$ and $\bar{\psi}_{l}$, and the other Wilson loop operators are identified with $\psi_{l,r}$. Aside from the Wilson loops and the identity operator, there are 15 twist operators, among which 7 have quantum dimension 1 and 8 have quantum dimension $n/2$. Their topological spin and fusion rule can be determined from the $S$ and $T$ matrices, similar to the odd $n$ case. In the untwisted sector (labeled by ++++), the $S$ and $T$ transformations constrained to the basis of states
@@
SIII Monte Carlo Simulation with Monopole Suppression
@@
Figure S2: Numerical phase diagram in the $(J_{v}/T,J_{s}/T)$ parameter space for various $\kappa$ obtained from Monte Carlo simulation for the gauge theory (S44). $|\mathbf{v}|$ and $s$ are the order parameters for the $U(1)$ and $\mathbb{Z}_{2}$ symmetries, respectively. The vertical dashed line in the upper left panel denotes the critical value for the 3D XY transition, $J_{v}/T=0.45422$, and the horizontal dashed line in the upper right panel denotes that for the 3D Ising transition, $J_{s}/T=0.22165$.
From Fig. S2 it can be seen that the multicritical point is always stable in the range of $\kappa$ we consider. A more rigorous argument about the stability of the multicritical point is given as follows. First consider a three-dimensional fixed volume $\tilde{\mathcal{V}}$ (preserved under the RG flow) as a subspace of the five-dimensional phase diagram of the effective field theory (12) in the main text. For such a $\tilde{\mathcal{V}}$ with $r_{v}$ and $r_{s}$ remaining irrelevant in it and containing the $O(2)\times O(1)$ WF fixed point, any RG trajectory—as long as it is not the direct trajectory from $O(2)\times O(1)$ to the $O(3)$ WF fixed point—will be driven away from the latter, since $t_{v}$ and $t_{s}$ are always relevant in the entire parameter space. This suggests the stability of the multicritical point within $\tilde{\mathcal{V}}$. On the other hand, the phase diagram $\mathcal{V}$ of the lattice model (11) in the main text spanned by $J_{v}$, $J_{s}$, and $\kappa$ is also an embedded three-dimensional subspace. Assuming that $r_{v}$ and $r_{s}$ are always irrelevant in $\mathcal{V}$, the lattice model phase diagram $\mathcal{V}$ can be deformed to the phase diagram in $\tilde{\mathcal{V}}$ with irrelevant $r_{v,s}$. This assumption is justified by the observation that upon varying $J_{s}$ and $J_{v}$ in the lattice model (11), the universality class of transitions between stable phases remains Ising or XY, where the $r_{v}$ and $r_{s}$ are indeed irrelevant. Thus, the multicritical point should also be stable in the lattice model.
SIV The Complex Lattice Model
@@
Figure S3: Mean-field phase diagram of the complex $O(2)$ theory Eq. (S45) with $\lambda=1.5$ and $\Lambda=0.2$ when all the possible gauge fields are deconfined. The gauge structure of each phase is shown explicitly. The 3D XY and the 3D Ising transitions are marked in red and yellow, respectively. The direct transition between the $O(2)$ Coulomb phase and the $D(\mathbb{Z}_{n})$ phase, marked in pink, could be in a coexistence of the 3D XY and 3D Ising classes.

### v2

@@ Appendix A End Matter
Appendix SA Data of the dihedral quantum double
@@
where the function $\mathrm{Cell}_{n}[x]$ subtracts (or adds) integer multiples of $n$ to take $x$ back to the interval ((n1)/2,(n1)/2](-(n-1)/2,(n-1)/2], and we denoted a generic element of $D_{n}$ by $g=C_{n}^{r}M^{s}$. The $R$-symbol is [77, 18]
@@
Appendix SB Equivalence between $S[O(2)\times O(2)]$ BF theory and $D_{n}$ quantum double
SB.1 Ground State Degeneracy and Quantum Dimension
@@
SB.2 Topological Spins and Fusion Rules
@@
Their fusion rule, consistent with those of $\psi$ anyons in $D^{\omega}(D_{n})$, is the same as Eq. (8). To be concrete, here $W_{p,0}$ and $W_{p,\frac{n}{2}}$ are identified as $\psi_{l}$ and $\bar{\psi}_{l}$, and the other Wilson loop operators are identified with $\psi_{l,r}$. Aside from the Wilson loops and the identity operator, there are 15 twist operators, among which 7 have quantum dimension 1 and 8 have quantum dimension $n/2$. Their topological spin and fusion rule can be determined from the $S$ and $T$ matrices, similar to the odd $n$ case. In the untwisted sector (labeled by ++++), the $S$ and $T$ transformations constrained to the basis of states
@@
Appendix SC Monte Carlo Simulation with Monopole Suppression
@@
A more rigorous argument about the stability of the multicritical point is given as follows. First consider a three-dimensional fixed volume $\tilde{\mathcal{V}}$ (preserved under the RG flow) as a subspace of the five-dimensional phase diagram of the effective field theory (12) in the main text. For such a $\tilde{\mathcal{V}}$ with $r_{v}$ and $r_{s}$ remaining irrelevant in it and containing the $O(2)\times O(1)$ symmetric fixed point, any RG trajectory—as long as it is not the direct trajectory from $O(2)\times O(1)$ to the $O(3)$ symmetric fixed point—will be driven away from the latter, since $t_{v}$ and $t_{s}$ are always relevant in the entire parameter space. This suggests the stability of the multicritical point within $\tilde{\mathcal{V}}$. On the other hand, the phase diagram $\mathcal{V}$ of the lattice model (11) in the main text spanned by $J_{v}$, $J_{s}$, and $\kappa$ is also an embedded three-dimensional subspace. Assuming that $r_{v}$ and $r_{s}$ are always irrelevant in $\mathcal{V}$, the lattice model phase diagram $\mathcal{V}$ can be deformed to the phase diagram in $\tilde{\mathcal{V}}$ with irrelevant $r_{v,s}$. This assumption is justified by the observation that upon varying $J_{s}$ and $J_{v}$ in the lattice model (11), the universality class of transitions between stable phases remains Ising or $U(1)$, where the $r_{v}$ and $r_{s}$ are indeed irrelevant. Thus, the multicritical point should also be stable in the lattice model.
Appendix SD The Complex Lattice Model
@@
Figure S2: Mean-field phase diagram of the complex $O(2)$ theory Eq. (S45) with $\lambda=1.5$ and $\Lambda=0.2$ when all the possible gauge fields are deconfined. The gauge structure of each phase is shown explicitly. The 3D $U(1)$ and the 3D Ising transitions are marked in red and yellow, respectively. The direct transition between the $O(2)$ Coulomb phase and the $D(\mathbb{Z}_{n})$ phase, marked in pink, could be in a coexistence of the 3D $U(1)$ and 3D Ising classes.